Grade 11Physical Chemistry

50 ml of a solution containing 1g each of na2co3,nahco3 and naoh was titrated with 1N hcl what will be the titre reading if only phenolphthalein is used as indicator

santosh
9 Years agoGrade 11

2 Answers

Vikas TU
9 Years ago
From mole 2 concept,
1/53 + 1/84 + 1/40 = 1 * V * 1
On solvin we get,
V = 55.77 ml
for valnce factor of Na2CO3  being 2 and 1 for remaining compunds.
varun
9 Years ago
HELLO SANTOSH,
first you need to do is 
equate the molarity or mole  equivalents of all the compounds present there,
Mequivalent of HCL =M equivalent of NaOH + M eqiuvalent of Na2CO3
                               =M equivalent of NaOH + M equivalent of NaHCO3
==> 1XV    = 1/40X 1000  + 1/106X 1000
                  =34ml
if you only phenophthalien indicator you will get the answer as 34ml
if you use methyl orange as an indicator you will get 55.77ml
THANK YOU